OpenStudy (anonymous):

A spherical tank contains 81.637 gallons of water at timet=0 minutes. For the next 6 minutes, water flows out ofthe tank at the rate of 9sin square root(t+1) gallons per minute. How many gallonsof water are in the tank at the end of the 6 minutes? why is it 36.606?

OpenStudy (bahrom7893):

Okay after 6 minutes: There is 9*Sin(sqrt(6+1)) amount of water leaked.

OpenStudy (bahrom7893):

So you have original: 81.637 - amount leaked = 81.637 - 9*Sin(sqrt(7))

OpenStudy (anonymous):

that does not equal 36.606

OpenStudy (bahrom7893):

It doesn't? Let me see

OpenStudy (anonymous):

its a related rates question

OpenStudy (bahrom7893):

Ohh

OpenStudy (bahrom7893):

this is so weird.. I've never encountered one like this.

OpenStudy (anonymous):

yeah ap calculus sucks

OpenStudy (bahrom7893):

I am pretty good at RR, but I don't remember seeing anything like this. V = (4/3)*pi*r^3 dV/dt = 4*pi*r^2(dr/dt)

OpenStudy (bahrom7893):

Water is flowing out so: dv/dt = -9sin[sqrt(t+1)]

OpenStudy (bahrom7893):

that gives me nothing. Sorry i'm lost. Let's see if others can help. @amistre64 , or @satellite73 or @KingGeorge . Can any of you guys take a look?

OpenStudy (kinggeorge):

Since you're given a rate, you should see that as a derivative. So \[{d \over dt}=9 \sin(\sqrt{t+1}) \]Just to make sure, this is for a calculus class correct?

OpenStudy (bahrom7893):

yea, i think. RR is calc

OpenStudy (bahrom7893):

So integrate that from 0 to 6?

OpenStudy (kinggeorge):

And that should have a negative in front. But yes, I would integrate that from 0 to 6 first.

OpenStudy (kinggeorge):

And then you add what you get to 81.637

OpenStudy (kinggeorge):

It's a rather nasty integration problem, given that wolfram seems to be doing two u-substitutions followed by an integration by parts.

OpenStudy (anonymous):

thx i'll try that

OpenStudy (kinggeorge):

Although it seems to me as if you could get by with only one u-sub then an integration by parts if you let \(u=\sqrt{t+1}\)

OpenStudy (kinggeorge):

In conclusion, you should be doing \[V= 81 + \int\limits_0^6 -9\sin(\sqrt{t+1}) \quad dt\]Where \(V\) is the amount of water left.